# A Practical Approach Towards Inertia Estimation Using Ambient Synchrophasor Data

### v1

@@ I Introduction
We first estimate the rotor speeds and acceleration from PMUs placed at a few of the network buses by leveraging the frequency divider formula [5], and then solve the resulting linear estimation problem in the presence of bounded uncertainties. Next, we determine the inertia constant ($H$) and damping constant ($D$) of individual generators from the swing equation while treating the mechanical power input ($p_{m}$) as a constant. However, since we do not know beforehand the length of the actual windows for which $p_{m}$ is constant, considerable fluctuations occur in the $H$ and $D$ estimates. We use the fluctuations in the estimates to figure out time intervals for which $p_{m}$ is an unknown constant, and then solve for all three unknown parameters ($H$, $D$, $p_{m}$) through a partitioned form of the swing equation. We draw insights from actual power system operations to ensure practicality of the proposed approach. Our main contributions are:
@@ II Proposed Approach for Rotor Speed and Acceleration Estimation
The frequency divider formula introduced in [5] establishes a linear relationship between rotor speeds of synchronous machines and the frequency of network buses. Combining this formula with the knowledge of network topology and bus frequencies obtained via PMUs, one can estimate the rotor speeds (and subsequently, acceleration), which can then be used to estimate inertia from the swing equation. However, network parameters are subject to changes and PMU measurements have errors, implying that the estimation problem will be subject to uncertainties. These aspects of the problem are discussed in this section, while the proposed approach for inertia estimation is described in the next section.
@@ II-A Frequency Divider and Spatial Correlation
Consider a power system having $N$ buses and $M$ point of interconnection (POI) buses to which generators are connected. For such a system, the frequency divider formula can be mathematically expressed in the following way [5]:
$0_{N}=B_{BG}\omega_{G}(t)+B_{BB}\omega_{B}(t),$ (1)
where $\omega_{B}(t)\in\mathbb{R}^{N}$ is the vector of bus frequencies; $\omega_{G}(t)\in\mathbb{R}^{M}$ is the vector of rotor speeds; and the matrices are:
@@
$\begin{bmatrix}0_{M}\\ 0_{N}\end{bmatrix}=\begin{bmatrix}\operatorname{diag}(b_{Gb})&B_{Bb}\\ 0_{N\times M}&B_{Bl}\end{bmatrix}\begin{bmatrix}\omega_{G}\\ \omega_{B}\end{bmatrix}$ (2)
@@
${\omega}_{G}=-\operatorname{diag}(b_{Gb})^{-1}B_{Bb}{\omega}_{B}$ (3)
@@
$\omega_{B,me}=E{\omega}_{B}$ (4)
where $\omega_{B,me}\in\mathbb{R}^{L\times 1}$ denotes the measured bus frequencies, and $E$ $\in\mathbb{R}^{L\times N}$ is an identity-like matrix whose rows corresponding to bus frequencies that are not measured by PMUs, have been removed. Next, by combining (4) with the second row of (2), the following relation between the generators’ rotor speeds and measured bus frequencies can be established [6]:
${\omega}_{G}$ =$=\left\{\operatorname{diag}(b_{Gb})^{-1}B_{Bb}\begin{bmatrix}E\\ B_{Bl}\end{bmatrix}^{+}\begin{bmatrix}I_{L}\\ 0_{N\times L}\end{bmatrix}\right\}\omega_{B,me}$ (5)
where, $I_{L}$ $\in\mathbb{R}^{L\times L}$ is the identity matrix, and $[\cdot]^{+}$ denotes the pseudo-inverse. Finally, by denoting the terms in curly brackets in (5) by $C\in\mathbb{R}^{M\times L}$, we establish the following linear relation between rotor speeds and measured bus frequencies:
${\omega}_{G}=C\omega_{B,me}$ (6)
To uniquely estimate all rotor speeds using (6), at a minimum, every POI bus frequency must be measured by PMUs. Going forward, we assume that is the case and set $L=M$. Note that one can show $C$ is full rank in this case. Finally, since PMUs give rate-of-change-of-frequency measurements and $C$ is time-invariant (as long as topology does not change), (6) can be extended to solve for rotor acceleration as well.
@@ II-B Estimation in Presence of Bounded Data Uncertainties
It has been documented in prior literature that network parameter uncertainties are bounded to within $\pm$30of their database values [7], while errors in the PMU-measured bus frequencies in the U.S. lie within $\pm$ 0.008 Hz [8]. To effectively incorporate knowledge of these bounds into our problem formulation, we treat the estimation problems as BDU problems. Taking (6) as an example of a BDU problem, we solve for $\hat{\omega}_{G}$ by performing the following $\operatorname{min-max}$ optimization:
$\min_{{\omega}_{G}}\max_{\begin{array}[]{c}\scriptstyle\|\delta\mathbb{A}\|_{2}\leq\eta\\[-4.0pt] \scriptstyle\|\delta{\omega}_{B,me}\|_{2}\leq\eta_{b}\\[-4.0pt] \end{array}}{\|(\mathbb{A}+\delta\mathbf{\mathbb{A}}){\omega}_{G}-({\omega}_{B,me}+\delta{\omega}_{B,me})\|_{2}}$
@@
      For $\eta\geq\tau_{2}$, the unique solution is $\hat{\omega}_{G}=0$.
@@
      For $\tau_{1}<\eta<\tau_{2}$, the unique solution is given by $\hat{\omega}_{G}=(\mathbb{A}^{\top}\mathbb{A}+\hat{\psi}I)^{-1}\mathbb{A}^{\top}{\omega_{B,me}}$, where $\hat{\psi}$ is the positive root of the secular equation, $\mathbb{G}(\hat{\psi})$ = 0.
@@
      For $\eta\leq\tau_{1}$, the unique solution is $\hat{\omega}_{G}=VS^{-1}w_{B,me}=\mathbb{A}^{+}{\omega_{B,me}}$.
@@
      For $\eta=\tau_{1}=\tau_{2}$, there are infinitely many solutions: $\hat{\omega}_{G}=\beta VS^{-1}w_{B,me}=\beta\mathbb{A}^{+}{\omega_{B,me}}$, for any $0\leq\beta\leq 1$.
@@ III-B A Partitioned Formulation for Inertia Estimation
Using the knowledge of when the estimates obtained from (9) have become inconsistent within the economic dispatch period, we partition the discrete-time interval $t=0,\ldots,T$ into $k$ non-overlapping intervals, denoted as $[t_{i-1}+d_{i-1},t_{i})$ for $i=1,\ldots,k$, where $t_{0}=0$, $d_{0}=0$, and $t_{k}=T$. Note that each of these intervals can have an arbitrary length. Now, since $p_{m}$ remains constant within each interval but may differ between intervals, we have $\bar{p}_{m}(t)=\bar{p}_{m_{i}}$ for $t\in[t_{i-1}+d_{i-1},t_{i})$. Finally, for $i\in\{1,\ldots,k\}$, it follows that:
@@ IV Simulation Results
We evaluate the proposed approach on the IEEE 14-bus and 39-bus systems. The former has no CIGs, but the latter has three CIGs at buses 2, 29, and 39. We generate data by creating random fluctuations in the loads by varying them within ±$\pm 0.1\%$ of their actual value [10]. Bounded perturbations specified in [7] and [8] are added to the network parameters and PMU measurements, respectively. The $p_{e}$ of the generators is changed every $8$ seconds, and the dynamic simulations are run for $40$ seconds in PSS/E. We use the absolute relative error ($\mathrm{ARE}$) metric for evaluating the performance.
@@ IV-A IEEE 14-Bus System
This system has five generators labeled {$G_{1}$, $G_{2}$, $G_{3}$, $G_{4}$, $G_{5}$}. The variation in $p_{e}$ and $p_{m}$ for $G_{1}$ is shown in Fig. 2. However, the knowledge of the interval after which the change occurs ($8$ seconds, in this case), is not known a priori. Therefore, we first solve (9) over the entire $40$-second time interval using a rolling window of size $0.33$ seconds. Figs. 3 and 4 demonstrate that the fluctuations in $p_{m}$ (shown in Fig. 2) introduce extremely high variability in the estimates of $H$ and $D$. Note that the regions of the plots of %$\%H_{\mathrm{ARE}}$ and %$\%D_{\mathrm{ARE}}$ which are blank (e.g., first $120$ rolling windows in Fig. 4) were due to the estimation algorithm not converging because of the ill-conditioning of the linear system of equations. One way to avoid this problem is by increasing the size of the rolling window. The optimal window size should be such that (i) the linear estimation problems derived in Section III are rank-sufficient, and (ii) there are multiple windows in which $p_{m}$ has the time to settle down and become flat.
@@
[…]
Now, by analyzing the variability in the estimates of $H$ and $D$, we identify time intervals where $p_{m}$ remains constant. Then, we use (12) to estimate $H$, $D$, and $p_{m}$ for those time intervals only. The results are shown in Table I for all five generators of the 14-bus system. Note that the entries in the last column indicate the average %$\%\mathrm{ARE}$ for $p_{m}$ across different time intervals. The very low values of the %$\%\mathrm{ARE}$ confirm the robustness of the proposed approach.
[…]
@@ IV-B IEEE 39-Bus System with Renewable Energy Resources
The modified IEEE 39-bus system consists of 3 renewable resources and 10 generators labeled {$G_{1}$, $\ldots$, $G_{10}$}. Estimation of $H$ and $D$ using (9) over the $40$-second interval results in Figs. 5 and 6. The figures confirm that the estimated values of $H$ and $D$ fluctuate considerably. Note that in contrast to Figs. 3 and 4 which showed the fluctuations in the %$\mathrm{ARE}$, Figs. 5 and 6 depict the variations in the actual estimates. However, the inferences drawn are the same (namely, that changes in $p_{m}$ severely deteriorate estimation performance). By examining these plots, intervals in which $p_{m}$ remains constant are identified and the remaining rolling windows are discarded. The %$\%\mathrm{ARE}$ of $H$, $D$, and $p_{m}$ estimates obtained using (12), after omitting the intervals where $p_{m}$ is varying, are shown in Table II. The %$\%\mathrm{ARE}$ values (particularly for $\hat{H}$) are even lower for this system confirming the accuracy and widespread applicability of the proposed approach.
@@
[…]
H ×$\times 10^{-4}$ D $p_{m,avg}$
@@ V Conclusion
This paper presents a novel approach for estimating three parameters—namely, inertia constant, damping constant, and mechanical power input—of every generator from ambient PMU data. The proposed approach relaxes the constant mechanical power input assumption and accounts for the presence of bounded uncertainties in the network parameters and PMU measurements. First, rotor speed and acceleration are estimated for all the generators from PMUs placed only at the POI buses by embedding frequency spatial correlations into the frequency divider formula. Then, these estimates are inserted into the swing equation to check for consistency of the inertia and damping estimates. Considerable variations in the resulting estimates are indicators of changes in mechanical power inputs. This knowledge is used to determine intervals for which the mechanical power input does not change. Finally, these intervals are fed into a partitioned form of the swing equation to estimate the three desired parameters as well as the system-level inertia constant. The results obtained using two IEEE test systems affirm the accuracy, robustness, and widespread applicability of the proposed approach. The focus of future research will be on investigating this approach in the presence of intermittency and variability typically encountered by CIGs in the field.

### v2

@@ I Introduction
We first estimate rotor speeds and acceleration from PMUs by extending the frequency divider formula (FDF) [5], and then solve the resulting linear estimation problem in the presence of bounded uncertainties. Next, we determine the inertia constant ($H$) and damping constant ($D$) of individual generators from the swing equation while treating the mechanical power input ($p_{m}$) as a constant. However, since we do not know beforehand the length of the actual windows for which $p_{m}$ is constant, considerable fluctuations occur in the $H$ and $D$ estimates. We use the fluctuations in the estimates to figure out time intervals for which $p_{m}$ is an unknown constant, and then solve for all three unknown parameters ($H$, $D$, $p_{m}$) through a partitioned form of the swing equation. We draw insights from actual power system operations to ensure practicality of the proposed approach. Our main contributions are:
@@ II Proposed Approach for Rotor Speed and Acceleration Estimation
The FDF establishes a linear relationship between rotor speeds of synchronous machines and frequency of network buses. Combining the FDF with the knowledge of network topology and bus frequencies obtained via PMUs, one can estimate the rotor speeds and acceleration, which can then be used to estimate inertia from the swing equation. However, network parameters are subject to changes and PMU measurements have errors, implying that the estimation problem will be subject to uncertainties. These aspects of the problem are discussed in this section, while the proposed approach for inertia estimation is described in the next section.
@@ II-A Frequency Divider and Spatial Correlation
Consider a power system having $N$ buses and $M$ point of interconnection (POI) buses to which generators are connected. For such a system, the FDF can be mathematically expressed in the following way [5]:
$0_{N}=B_{BG}(\omega_{G}(t)-\mathbf{1})+B_{BB}(\omega_{B}(t)-\mathbf{1}),$ (1)
where $\omega_{B}(t)\in\mathbb{R}^{N}$ is the vector of bus frequencies; $\omega_{G}(t)\in\mathbb{R}^{M}$ is the vector of rotor speeds; $\mathbf{1}$ is an appropriately sized vector indicating nominal frequency (angular/bus) in p.u.; and the matrices are:
@@
$\begin{bmatrix}0_{M}\\ 0_{N}\end{bmatrix}=\begin{bmatrix}\operatorname{diag}(b_{Gb})&B_{Bb}\\ 0_{N\times M}&B_{Bl}\end{bmatrix}\begin{bmatrix}\omega_{G}-\mathbf{1}\\ \omega_{B}-\mathbf{1}\end{bmatrix}$ (2)
@@
$({\omega}_{G}-\mathbf{1})=-\operatorname{diag}(b_{Gb})^{-1}B_{Bb}({\omega}_{B}-\mathbf{1})$ (3)
@@
$\omega_{B,me}=E({\omega}_{B}-\mathbf{1})$ (4)
where $\omega_{B,me}\in\mathbb{R}^{L\times 1}$ denotes the measured bus frequency deviations, and $E$ $\in\mathbb{R}^{L\times N}$ is an identity-like matrix whose rows corresponding to bus frequencies that are not measured by PMUs, have been removed. Next, by combining (4) with the second row of (2), the following relation between the generators’ rotor speed deviations and the measured bus frequency deviations can be established [6]:
$({\omega}_{G}-\mathbf{1})$ $=\left\{\operatorname{diag}(b_{Gb})^{-1}B_{Bb}\begin{bmatrix}E\\ B_{Bl}\end{bmatrix}^{+}\begin{bmatrix}I_{L}\\ 0_{N\times L}\end{bmatrix}\right\}\omega_{B,me}$ (5)
where, $I_{L}$ $\in\mathbb{R}^{L\times L}$ is an identity matrix, and $\cdot$$^{+}$ denotes pseudo-inverse. Finally, by denoting the terms in curly brackets in (5) by $C\in\mathbb{R}^{M\times L}$, we establish the following relation between rotor speed and measured bus frequency deviations:
$({\omega}_{G}-\mathbf{1})=C\omega_{B,me}$ (6)
To uniquely estimate all rotor speeds using (6), at a minimum, every POI bus frequency must be measured by PMUs. Going forward, we assume that is the case and set $L=M$. Note that one can show $C$ is full rank in this case. Finally, (6) can be used to solve for rotor acceleration as well.
@@ II-B Estimation in Presence of Bounded Data Uncertainties
It has been documented in prior literature that network parameter uncertainties are bounded to within $\pm$30% of their database values [7], while errors in PMU-measured bus frequencies in the U.S. lie within $\pm$ 0.008 Hz [8]. To incorporate knowledge of these bounds into our problem formulation, we treat (6) as a BDU problem and solve for $\hat{\omega}_{G}$ by performing the following $\operatorname{min-max}$ optimization:
$\min_{\omega_{G}}\max_{\begin{subarray}{c}\|\delta\mathbb{A}\|_{2}\leq\eta\\ \|\delta\omega_{B,me}\|_{2}\leq\eta_{b}\end{subarray}}\big{\|}(\mathbb{A}+\delta\mathbb{A})(\omega_{G}-\mathbf{1})-(\omega_{B,me}+\delta\omega_{B,me})\big{\|}_{2}$
@@
      For $\eta\geq\tau_{2}$, the unique solution is $\hat{\omega}_{G}=\mathbf{1}$.
@@
[…]
@@
      For $\eta\leq\tau_{1}$, the unique solution is $\hat{\omega}_{G}=VS^{-1}w_{B,me}+\mathbf{1}=\mathbb{A}^{+}{\omega_{B,me}}+\mathbf{1}$.
@@
      For $\eta=\tau_{1}=\tau_{2}$, there are infinitely many solutions: $\hat{\omega}_{G}=\beta VS^{-1}w_{B,me}+\mathbf{1}=\beta\mathbb{A}^{+}{\omega_{B,me}}+\mathbf{1}$, for any $0\leq\beta\leq 1$.
@@ III-B A Partitioned Formulation for Inertia Estimation
Using the knowledge of when the estimates obtained from (9) have become inconsistent within the economic dispatch period, we partition the discrete-time interval $t=0,\ldots,T$ into $k$ non-overlapping intervals, denoted as $t_{i-1}+d_{i-1},t_{i})$ for $i=1,\ldots,k$, where $t_{0}=0$, $d_{0}=0$, and $t_{k}=T$. Note that each of these intervals can have an arbitrary length. Now, since $p_{m}$ remains constant within each interval but may differ between intervals, we have $\bar{p}_{m}(t)=\bar{p}_{m_{i}}$ for $t\in[t_{i-1}+d_{i-1},t_{i})$. Finally, for $i\in\{1,\ldots,k\}$, it follows that:
@@ IV Simulation Results
We evaluate the proposed approach on the IEEE 14-bus and 39-bus systems. The former has no CIGs, but the latter has three CIGs at buses 2, 29, and 39. We generate data by creating random fluctuations in the loads by varying them within $\pm 0.1\%$ of their actual value [10]. Bounded perturbations specified in [7] and [8] are added to the network parameters and PMU measurements, respectively. The $p_{e}$ of the generators is changed every $8$ seconds, and the dynamic simulations are run for $40$ seconds in PSS/E. We use the absolute relative error ($\mathrm{ARE}$) metric for evaluating the performance.
@@ IV-A IEEE 14-Bus System
This system has five generators labeled $G_{1}$, $G_{2}$, $G_{3}$, $G_{4}$, $G_{5}$}. The variation in $p_{e}$ and $p_{m}$ for $G_{1}$ is shown in Fig. 2. However, the knowledge of the interval after which the change occurs ($8$ seconds, in this case), is not known a priori. Therefore, we first solve (9) over the entire $40$-second time interval using a rolling window of size $0.33$ seconds. Figs. 3 and 4 demonstrate that the fluctuations in $p_{m}$ (shown in Fig. 2) introduce extremely high variability in the estimates of $H$ and $D$. Note that the regions of the plots of $\%H_{\mathrm{ARE}}$ and $\%D_{\mathrm{ARE}}$ which are blank (e.g., first $120$ rolling windows in Fig. 4) were due to the estimation algorithm not converging because of the ill-conditioning of the linear system of equations. One way to avoid this problem is by increasing the size of the rolling window. The optimal window size should be such that (i) the linear estimation problems derived in Section III are rank-sufficient, and (ii) there are multiple windows in which $p_{m}$ has the time to settle down and become flat.
@@
Figure 3: $H_{\mathrm{ARE}}$ (in $\%$) for $G_{1}$ in presence of variable mechanical power input
Figure 4: $D_{\mathrm{ARE}}$ (in $\%$) for $G_{1}$ for variable mechanical power input
Now, by analyzing the variability in the estimates of $H$ and $D$, we identify time intervals where $p_{m}$ remains constant. Then, we use (12) to estimate $H$, $D$, and $p_{m}$ for those time intervals only. The results are shown in Table I for all five generators of the 14-bus system. Note that the entries in the last column indicate the average $\%\mathrm{ARE}$ for $p_{m}$ across different time intervals. The very low values of the $\%\mathrm{ARE}$ confirm the robustness of the proposed approach.
TABLE I: IEEE 14-bus system: $\mathrm{ARE}$ (in $\%$) for $H$, $D$, and $p_{m}$ using partitioned form of swing equation
Generator $\mathrm{ARE}$ in $\%$
@@ IV-B IEEE 39-Bus System with Renewable Energy Resources
The modified IEEE 39-bus system consists of 3 renewable resources and 10 generators labeled $G_{1}$, $\ldots$, $G_{10}$}. Estimation of $H$ and $D$ using (9) over the $40$-second interval results in Figs. 5 and 6. The figures confirm that the estimated values of $H$ and $D$ fluctuate considerably. Note that in contrast to Figs. 3 and 4 which showed the fluctuations in the $\mathrm{ARE}$, Figs. 5 and 6 depict the variations in the actual estimates. However, the inferences drawn are the same (namely, that changes in $p_{m}$ severely deteriorate estimation performance). By examining these plots, intervals in which $p_{m}$ remains constant are identified and the remaining rolling windows are discarded. The $\%\mathrm{ARE}$ of $H$, $D$, and $p_{m}$ estimates obtained using (12), after omitting the intervals where $p_{m}$ is varying, are shown in Table II. The $\%\mathrm{ARE}$ values (particularly for $\hat{H}$) are even lower for this system confirming the accuracy and widespread applicability of the proposed approach.
@@
TABLE II: IEEE 39-bus system: $\mathrm{ARE}$ (in $\%$) for $H$, $D$, and $p_{m}$ using partitioned form of swing equation
Generator $\mathrm{ARE}$ in $\%$
H $\times 10^{-4}$ D $p_{m,avg}$
@@ V Conclusion
This paper presents a novel approach for estimating three parameters—namely, inertia constant, damping constant, and mechanical power input—of every generator from ambient PMU data. The proposed approach relaxes the constant mechanical power input assumption and accounts for the presence of bounded uncertainties in the network parameters and PMU measurements. First, rotor speed and acceleration are estimated for all the generators from PMUs placed only at the POI buses by embedding frequency spatial correlations into the FDF. Then, these estimates are inserted into the swing equation to check for consistency of the inertia and damping estimates. Considerable variations in the resulting estimates are indicators of changes in mechanical power inputs. This knowledge is used to determine intervals for which the mechanical power input does not change. Finally, these intervals are fed into a partitioned form of the swing equation to estimate the three desired parameters as well as the system-level inertia constant. The results obtained using two IEEE test systems affirm the accuracy, robustness, and widespread applicability of the proposed approach. The focus of future research will be on investigating this approach in the presence of intermittency and variability typically encountered by CIGs in the field.